# thanks: Part of the Instituto Avanzado de Cosmología (IAC) Collaboration

### v1

@@ Abstract
Observations with the Gaia satellite have confirmed that the satellite galaxies of the Milky Way are not distributed as homogeneously as expected. The same occurs in galaxies such as Andromeda and Centaurus A, where satellite galaxies around their host galaxies have been observed to have orbits aligned perpendicular to the galactic plane of the host galaxy. This problem is known for the Milky Way as Vast Polar Structure (VPOS). The Scalar Field Dark Matter Field (SFDM), also known as Ultralight-, Fuzzy-, BEC-, Axion-dark matter, proposes that dark matter is a scalar field, which in the non-relativistic limit follows the Schrödinger equation coupled to the Poisson equation. Although the scalar field here is classical, the Schrödinger equation contains a ground and excited states as part of its nature. In this work, we show that such quantum character of the SFDM can naturally explain the VPOS observed in galaxies. By taking into account the finite temperature corrections for a complex, self-interacting SF at very early epochs of the Universe, we show that the resulting ground and first excited states only, we can fit the rotation curves of the galaxies in a very simple way, and with the best-fit parameters obtained, we can explain the VPOS. We do this with particular galaxies, such as the Milky Way, Andromeda, Centaurus A, and other six galaxies whose satellites have been observed. From this result, it follows that the multistate SFDM is not distributed homogeneously around the galaxy, and therefore might explain the anisotropic distribution of the satellite galaxies. According to this result, this could be a general characteristic of the galaxies in the Universe. Finally, we also show how the scale of each galaxy depends on a parameter that is determined by the final temperature of the scalar field of the galaxy halo under study. This explains why different galaxies with SFDM give different values of the mass of the scalar field.
galaxy halos – Bose condensates
I Introduction
@@
Dark Matter (DM) stands as one of the most important scientific puzzles of the current century. Despite advances, over 95% of the Universe’s matter-energy content eludes comprehension, with DM comprising more than a quarter of this total. DM is the responsible of the formation of diverse structures in the Universe, ranging from superclusters of galaxies to dwarf galaxies scattered across the cosmos. Since its discovery in the early 1930s, unraveling the composition of dark matter has remained a significant challenge, representing one of the foremost enigmas in science.
@@
The most widely accepted and studied model that explains various cosmological observations, as well as those on large galaxies and galaxy cluster scales, is the cold dark matter (CDM). However, CDM presents some problems at galactic scales, e.g., the so-called “core-cusp” problem, the “missing satellites” problem, and all CDM models predict that satellite galaxies should move in homogeneous orbits within their host galaxies and be uniformly distributed. However, observations from current telescopes have revealed the satellite galaxies of three galaxies, the Milky Way (MW) (Pawlowski and Kroupa, 2013, 2020), Andromeda (M31) (Ibata et al., 2013; Conn et al., 2013), and Centaurus A (Cen A) (Müller et al., 2018, 2021), are not evenly distributed but exhibit a phenomenon known as Vast Polar Structure (VPOS), named in this way for the MW. Thus, the CDM paradigm faces challenges in its predictions, particularly on galactic scales (Weinberg et al., 2015). This has motivated the search for alternative models consistent with observational data (see e.g. Oks, 2021).
@@
One alternative is the Scalar Field Dark Matter (SFDM) model (see e.g. Matos and Urena-Lopez, 2001; Hui et al., 2017; Suárez et al., 2014), proposed first in (Ji and Sin, 1994; Lee and Koh, 1996), and systematically studied the next years by many groups (see e.g. Matos and Guzman, 2000; Urena-Lopez and Liddle, 2002; Matos et al., 2023). The SFDM model proposes that DM is a scalar field (SF), a spin-0 particle, governed by the Klein-Gordon equations driving the dynamics of the Universe (see e.g. Matos et al., 2023; Hui et al., 2017; Bernal et al., 2018, and references therein). The SFDM model is known by different names, including Fuzzy (Hu et al., 2000), Bose-Einstein (Boehmer and Harko, 2007; Rindler-Daller and Shapiro, 2010), Wave DM (Bray, 2010; Schive et al., 2014), among others. SFDM has gained popularity as one of the favored candidates to explain DM.
@@
This hypothesis can explain the observed rotation curves of stars and gas around galaxies (see e.g. Matos and Guzman, 2000; Bernal et al., 2018; Solís-López et al., 2024; Alvarez-Rios et al., 2024, among others). Also, the SFDM model can explain all the cosmological observations as well as the CDM model; it reproduces the observed mass power spectrum and the angular power spectrum (see e.g. Matos and Urena-Lopez, 2001; Schive et al., 2014; Hlozek et al., 2015, and references therein). The difference with CDM is that, given its quantum nature, the SFDM model presents “core” density profiles instead “cusp” ones as the CDM, and SFDM presents a natural cut-off in the mass power spectrum, implying that the theoretical number of satellite galaxies matches the observed quantities, solving in this way the “missing satellites” problem.
@@
Diverse proposals for the origin and nature of SFDM highlight the ongoing exploration and research efforts to better comprehend the mysterious nature of DM and its potential connections to fundamental physics. For instance, SFDM can arise from superstring theory (Hui et al., 2017); in this scenario, the SF is real and electrically neutral. Another possibility is to incorporate the SFDM Lagrangian into the Standard Model (SM) of particle physics, which would establish a connection between DM and the known particle physics of the Universe. (Rindler-Daller and Shapiro, 2014), a complex scalar field $\Phi$ is considered as a potential candidate for SFDM.
@@
The SFDM model has also faced some problems that have been solved. The first was discovered in Guzman and Urena-Lopez (2003), revealing the potential instability of galaxies under this model. The second issue concerned the possibility of having supermassive black holes at the center of galaxies consuming the entire SF halo. The latter problem was addressed showing that supermassive black holes can coexist with the SFDM halo (Urena-Lopez and Liddle, 2002; Barranco et al., 2011; Avilez et al., 2018; Padilla et al., 2021).
@@
The first problem was addressed by considering the quantum character of the SF taking into account multiple wave-functions as states or excited states of the system. In reference (Urena-Lopez and Bernal, 2010), the properties of gravitationally bound multistate configurations composed of spin-zero bosons in the Newtonian regime were studied. They found that the system remains stable when containing multiple states simultaneously. This result was corroborated in (Guzmán and Ureña López, 2020) and more recently in (Guzman, 2022). This finding is particularly important for the present work. The multistate solutions are also known as $l$-boson stars Alcubierre et al. (2018).
@@
Additionally, the multistate SFDM has been also addressed using quantum field theory at finite temperature in (Matos and Suárez, 2014; Robles and Matos, 2013a, b), for a real SF. This is the alternative we choose to explore in the present work, but for a complex SF. Our hypothesis involves a complex SF immersed in a thermal bath at finite temperature with the other SM particles, at the very early Universe. The SF being governed by a quartic SFDM potential, going through a spontaneous symmetry breaking (SSB) or phase transition, allowing the formation of SFDM halos with particles in a Bose-Einstein condensate ground state and excited states also present, treating them as atom-like systems. We obtain an effective SF mass that depends on the temperature of the thermal bath and the critical temperature of the halo at the moment the system collapsed.
@@
We assume that a complex SF is the DM of the Universe and is the dominant component contributing to the gravitational field. The electromagnetic fields are assumed to be small and have a negligible contribution to the overall gravity of the system. In a galaxy context, the SFDM can be effectively described by the non-relativistic limit of the Klein-Gordon (KG) equations, leading to the Schrödinger equation. By applying this limit to the Einstein-KG field equations, they reduce to the Schrödinger-Poisson (SP) system, which describes the behavior of SFDM halos in a non-relativistic, quantum mechanical framework.
@@
In this frame, we can interpret the SFDM halos of galaxies as macroscopic gravitational atoms with various quantum states similar to those observed in atomic physics, such as $s$, $p$, $d$, etc. The $s$ states correspond to spherical configurations, while the $p$ states resemble lobes or bubbles at the north and south poles of the galactic atom. Numerical simulations and evolution of the SP system performed in Guzmán and Ureña López (2020) have shown that the SF density profile exhibits a resemblance to the Legendre functions $P_{k}^{j}(\cos(\theta))$ of the hydrogen atom, where $k$ and $j$ are quantum numbers. This similarity indicates that the SFDM profile behaves similarly to a hydrogen atom, supporting the analogy between galaxies and atoms.
@@
This quantum behavior of the SFDM model is the main difference between the SFDM paradigm and other DM models. This feature is just the main point that allows us to explain various DM observations, especially at galactic scales, in particular the anisotropic distribution of satellite galaxies observed in the Milky Way, known as Vast Polar Structure (VPOS) (Pawlowski and Kroupa, 2013, 2020), Andromeda (Ibata et al., 2013; Conn et al., 2013), and Centaurus A (Müller et al., 2018, 2021), which is not explained by conventional CDM models. Thus, taking into account the quantum behavior of the SFDM, we can explain the behavior of the VPOS in a very simple and natural way, as proposed first in (Solís-López et al., 2021; Castro-Lara, 2021), which is the aim of this work.
@@
To do this, the article is organized as follows. In Section II, we provide a summary of the finite temperature SFDM model, presenting in II.1 the previous work done for a real SF, the successes of this model, and unsolved issues. In Subsection II.2, we start with the quartic potential for a complex SF, immersed in a thermal bath with the other components of the SM at very early stages of the Universe, and show the conditions of the SSB as the Universe cools down. In this section, we obtain the effective mass of the SF, which depends not only on the intrinsic mass of the field and its self-interaction parameter but also on the temperature of the halo at the moment the system collapses gravitationally and on the critical temperature at which the SSB occurs. In II.3, we present the evolution equations of the complex SF after the SSB takes place in a FLRW Universe, separating the SF function into its background part and a linear perturbation, and show the Newtonian limit to find the analytic solutions for the density, as we are interested in already formed galactic halos. In II.4 we perform a polar decomposition of the SF function into its norm and its phase. This separation allows us to rewrite the KG equations as a continuity and a Bernoulli equations. This is the hydrodynamic representation of the KG equation. This form of the KG equations allows us to point out that the only difference between a classical hydrodynamic system and the KG equation is a term known as the quantum potential. If such quantum term is zero, the system is classical. Therefore, we can say that the quantum character of the KG equation is contained in this term. In Subsection II.5 we find the exact solutions for the ground and first excited states through separation of variables, spherical harmonics, and Bessel polynomials, and we write the corresponding densities for these states. The main result shown here is that the final collapse of the SFDM is a halo in the form of an atom, with a first excited state with the form of two lobes along the north-south direction. To show if these resulting halos can explain the anisotropic distribution of the satellites in the MW, M31, CenA and another six MW-like galaxies, we fit the exact mSFDM solutions to the observed rotation curves of the galaxies and used the best-fit parameters to compare the resulting atom-like structures with the gravitational potential needed to explain the distribution of the satellites around the galaxies. In Section III we describe the observations we used, and the bulge and disc model we use in some cases. In Section IV we show the results of the fittings and the resulting structures of the two-lobes halos. Finally in section V we give some conclusions.
II Finite Temperature SFDM Model
@@
At the beginning of the Universe, we assume that SFDM was in thermal equilibrium with the other components of the Standard Model (SM) of particles but decoupled from them very early in the Universe’s history. Starting with the SF quartic potential $V$ coupled to a thermal bath with temperature $T$, we have (Kolb and Turner, 1990; Dalfovo et al., 1999)
$V=-m^{2}\Phi\Phi^{*}+\frac{\lambda}{2}(\Phi\Phi^{*})^{2}+\frac{\lambda}{4}\Phi\Phi^{*}T^{2}+\frac{\pi^{2}}{90}T^{4},$ (1)
@@
for the case when $T\gg m$, where $m$ is the mass of the SF and $\lambda$ the self-interaction parameter between SF particles. Observations indicate that the self-interaction strength $\lambda$ must be extremely small but non-zero, as constrained by nucleosynthesis (Li et al., 2014). The mass $m$ of the SFDM is considered ultralight, which implies that any perturbation in its temperature is comparable to this mass in energy.
@@
Now, of particular interest for cosmology is the theoretical expectation that at high temperatures, symmetries that are spontaneously broken today were restored and during the evolution of the Universe, there were phase transitions associated with the spontaneous breakdown of gauge symmetries (Kolb and Turner, 1990). So we expect that, as the Universe expanded, the SFDM particles and the other SM components, cooled down and underwent a phase transition at a critical temperature $T_{c}$ very early in the cosmic time. At this point, occurred a spontaneous symmetry breaking (SSB) of the SF potential.
@@
According to the standard cosmological model, after inflation, quantum fluctuations evolved into classical fluctuations, leading to the collapse of the SFDM and the formation of cosmic structures (Matos and Urena-Lopez, 2001; Urena-Lopez and Bernal, 2010; Hlozek et al., 2015). As the Universe expanded, the volume of these fluctuations increased, causing the SF to cool and form Bose-Einstein condensates. These condensates subsequently formed galaxy halos, where most SFDM particles settled into the ground state. However, gravitational forces caused the fluctuations to collapse again, decreasing their volume and increasing the halo’s temperature, resulting in some SFDM particles transitioning to excited states.
@@
In reference Guzmán and Ureña López (2020), it was demonstrated that if the majority of SFDM particles remain in the ground state while a portion of them occupy at least one excited state, the system stabilizes and remains stable for at least the age of the Universe, if the mass of the SFDM particle is sufficiently small, about $m\approx 10^{-22}\text{eV}/c^{2}$. This stability mechanism, involving a mixture of ground and excited states for SFDM particles within halos, is crucial for ensuring the formation and persistence of cosmic structures.
II.1 Previous Work and Unsolved Issues

### v3

# A Natural Explanation of the VPOS from Multistate Scalar Field Dark Matter

@@ Abstract
Observations with the Gaia satellite have confirmed that the satellite galaxies of the Milky Way are not distributed as homogeneously as expected. The same occurs in galaxies such as Andromeda and Centaurus A, where satellites around their host galaxies have been observed to have orbits aligned perpendicular to the galactic plane of the host galaxy. This problem is known for the Milky Way as Vast Polar Structure (VPOS). The Scalar Field Dark Matter Field (SFDM), also known as Ultralight-, Fuzzy-, BEC-, and Axion-dark matter, proposes dark matter is a scalar field, which in the non-relativistic limit follows the Schrödinger equation coupled to the Poisson equation. Although the SF here is classical, the Schrödinger equation contains a ground and excited states as part of its nature. In this work, we show that such quantum character of the SFDM can naturally explain the VPOS observed in galaxies. By taking into account the finite temperature corrections for a complex, self-interacting SF at very early epochs of the Universe, we show that with the ground and first excited states in the Newtonian limit, we can fit the rotation curves of the host galaxies. With the best-fit parameters obtained, we can explain the VPOS. We do this with particular galaxies, such as the Milky Way, Andromeda, Centaurus A, and 6 other galaxies whose satellites have been observed. This result shows that the multistate SFDM is not distributed homogeneously around the galaxy, and therefore might explain the anisotropic distribution of the satellite galaxies. According to this result, this could be a general characteristic of the galaxies in the Universe. Finally, we also show how the scale of each galaxy depends on a parameter determined by the final temperature of the SF galactic halo under study. This might explain why different galaxies with SFDM give different values of the mass of the SF.
1 Introduction
@@
Dark Matter (DM) stands as one of the most important scientific puzzles of the current century. Despite advances, over 95% of the Universe’s matter-energy content eludes comprehension, with DM comprising more than a quarter of this total. DM is responsible for the formation of diverse structures in the Universe, ranging from superclusters of galaxies to dwarf galaxies scattered across the cosmos. Since its discovery in the early 1930s, unraveling the composition of dark matter has remained a significant challenge, representing one of the foremost enigmas in science.
@@
The most widely accepted and studied model that explains various cosmological observations, as well as those on large galaxies and galaxy cluster scales, is cold dark matter (CDM). However, CDM presents some problems at galactic scales, for example the so-called “core-cusp” problem, the “missing satellites” problem, and all CDM models predict that satellite galaxies should move in homogeneous orbits within their host galaxies and be uniformly distributed [see e.g. 1, 2, 3, 4, 5]. However, observations from current telescopes have revealed the satellite galaxies of three galaxies, the Milky Way (MW) [6, 7], Andromeda (M31) [8, 9], and Centaurus A (Cen A) [10, 11], are not evenly distributed but exhibit a phenomenon known as Vast Polar Structure (VPOS), named in this way for the MW. Although the CDM model appears to be transiently compatible with this observation [12], it is challenging to explain this phenomenon in multiple galaxies, especially the coincidence that these three galaxies in our neighborhood exhibit this phenomenon. Thus, the CDM paradigm faces challenges in its predictions, particularly on galactic scales [13]. This has motivated the search for alternative models consistent with observational data [see e.g. 14].
@@
One alternative is the Scalar Field Dark Matter (SFDM) model [see e.g. 15, 16, 17], proposed first in [18, 19], and systematically studied the next years by many groups [see e.g. 20, 21, 22]. The SFDM model proposes that DM is a scalar field (SF), a spin-0 particle, governed by the Klein-Gordon equations driving the dynamics of the Universe [see e.g. 22, 16, 23, and references therein]. The SFDM model is known by different names, including Fuzzy [24], Bose-Einstein [25, 26], Wave DM [27, 28], among others. SFDM has gained popularity as one of the favored candidates to explain DM.
@@
This hypothesis can explain the observed rotation curves of stars and gas around galaxies [see e.g. 20, 23, 29, 30, among others]. Also, the SFDM model can explain all the cosmological observations as well as the CDM model; it reproduces the observed mass power spectrum and the angular power spectrum [see e.g. 15, 28, 31, and references therein]. The difference with CDM is that, given its quantum nature, the SFDM model presents “core” density profiles instead “cusp” ones, and SFDM presents a natural cut-off in the mass power spectrum, implying that the theoretical number of satellite galaxies matches the observed quantities, solving in this way the “missing satellites” problem.
@@
Diverse proposals for the origin and nature of SFDM highlight the ongoing exploration and research efforts to comprehend better the mysterious nature of DM and its potential connections to fundamental physics. For instance, SFDM can arise from superstring theory [16]; in this scenario, the SF is real and electrically neutral. Another possibility is to incorporate the SFDM Lagrangian into the Standard Model (SM) of particle physics, establishing a connection between DM and the known particle physics of the Universe. In [32], a complex scalar field $\Phi$ is considered a potential candidate for SFDM. In [33], the authors examine the possibility that SFDM consists of charged dark bosons whose gauge charge is associated with a dark photon.
@@
The SFDM model has also faced some problems that have been solved. The first was discovered in [34], revealing the potential instability of galaxies under this model. The second issue concerned the possibility of having supermassive black holes at the center of galaxies consuming the entire SF halo. The latter problem was addressed showing that supermassive black holes can coexist with the SFDM halo [21, 35, 36, 37].
@@
The first problem was addressed by considering the quantum character of the SF taking into account multiple wave functions as states or excited states of the system. In reference [38], the properties of gravitationally bound multistate configurations composed of spin-zero bosons in the Newtonian regime were studied. They found that the system remains stable when containing multiple states simultaneously. This result was corroborated in [39] and more recently in [40]. This finding is particularly important for the present work. The multistate solutions are also known as $l$-boson stars [41].
@@
The multistate SFDM has also been studied using quantum field theory at finite temperature in [42, 43, 44], for a real SF. This is the alternative that we choose to explore in the present work, but for a complex SF. Our hypothesis involves a complex SF immersed in a thermal bath at finite temperature with the other SM particles, at the very early Universe. The SF is governed by a quartic SFDM potential, going through an spontaneous symmetry breaking (SSB) or phase transition, allowing the formation of SFDM halos with particles in a Bose-Einstein condensate ground state, and excited states also present, treating them as atom-like systems. We obtain an effective SF mass that depends on the temperature of the thermal bath and the critical temperature of the halo at the moment the system collapses.
@@
We assume that a complex SF is the DM of the Universe and is the dominant component contributing to the gravitational field. The electromagnetic fields are assumed to be small and have a negligible contribution to the overall gravity of the system. In a galaxy context, the SFDM can be effectively described by the non-relativistic limit of the Klein-Gordon (KG) equations. By applying this limit to the Einstein-KG field equations, they reduce to the Schrödinger-Poisson (SP) system, which describes the behavior of SFDM halos in a non-relativistic, quantum mechanical framework.
@@
In this frame, we can interpret the SFDM halos of galaxies as macroscopic gravitational atoms with various quantum states similar to those observed in atomic physics, such as $s$, $p$, $d$, etc. The $s$ states correspond to spherical configurations, while the $p$ states resemble lobes or bubbles at the north and south poles of the galactic atom. Numerical simulations and evolution of the SP system performed in [39] have shown that the SF density profile exhibits a resemblance to the Legendre functions $P_{k}^{j}(\cos\theta)$ of the hydrogen atom, where $k$ and $j$ are quantum numbers. This similarity indicates that the SFDM profile behaves similarly to a hydrogen atom, supporting the analogy between galaxies and atoms.
@@
This quantum behavior of the SFDM model is the main difference between the SFDM paradigm and other DM models. This feature is just the main point that allows us to explain various DM observations, especially at galactic scales, in particular the anisotropic distribution of satellite galaxies observed in the Milky Way, known as Vast Polar Structure (VPOS) [6, 7], Andromeda [8, 9], and Centaurus A [10, 11], which conventional CDM models do not explain. Thus, taking into account the quantum behavior of the SFDM, we can explain the behavior of the VPOS in a very simple and natural way, as proposed first in [45, 46], which is the aim of this work [see also 47].
@@
To do this, the article is organized as follows. In Section 2, we provide a summary of the finite temperature SFDM model, presenting in 2.1 the previous work done for a real SF, the successes of such model, and unsolved issues. In subsection 2.2, we start with the quartic potential for a complex SF, immersed in a thermal bath with the other components of the SM at very early stages of the Universe, and show the conditions of an SSB as the Universe cools down. In this subsection, we obtain the effective mass of the SF, which depends not only on the intrinsic mass of the field and its self-interaction parameter but also on the temperature of the halo at the moment the system gravitationally collapses and on the critical temperature at which the SSB occurs. In 2.3, we present the evolution equations of the complex SF after the SSB takes place in an FLRW Universe, separating the SF function into its background part and a linear perturbation, and show the Newtonian limit to find the analytic solutions for the density, as we are interested in already formed galactic halos. In subsection 2.4, we perform a polar decomposition of the SF function into its norm and its phase. This separation allows us to rewrite the KG equations as a continuity and Bernoulli equations. This is the hydrodynamic representation of the KG equations. This form of the KG equations allows us to point out that the only difference between a classical hydrodynamical system and the KG equation is a term known as the quantum potential. If such a quantum term is zero, the system is classical. Therefore, we can say that the quantum character of the KG equation is contained in this term. In subsection 2.5 we find exact solutions for the ground and first excited states through the separation of variables, spherical harmonics, and Bessel polynomials, and we write the corresponding densities of those states. The main result shown here is that the final collapse of the SFDM is a halo in the form of an atom, with a first excited state in the form of two lobes along the north-south direction. To show if these resulting halos can explain the anisotropic distribution of the satellites in the MW, M31, CenA, and another 6 MW-like galaxies, we fit the exact mSFDM solutions to the observed rotation curves of the galaxies and used the best-fit parameters to compare the resulting atom-like structures with the gravitational potential needed to explain the distribution of the satellites around the galaxies. In Section 3, we describe the bulge and disc models we use to fit the rotation curves of the galaxies. In Section 4, we explain the MCMC calibration method we use to fit the models with the observations. In Section 5 we describe the observations we used, and the priors for the calibration method. In Section 6 we show the results of the fittings and the resulting structures of the two-lobe halos. Finally, in Section 7 we present some conclusions.
2 Finite Temperature SFDM Model
@@
At the beginning of the Universe, we assume that SFDM was in thermal equilibrium with the other components of the Standard Model (SM) of particles but decoupled from them very early in the Universe’s history. Starting with the SF quartic potential $V$ coupled to a thermal bath with temperature $T$, we have [48, 49]
$V=-m^{2}\Phi\Phi^{*}+\frac{\lambda}{2}(\Phi\Phi^{*})^{2}+\frac{\lambda}{4}\Phi\Phi^{*}T^{2}+\frac{\pi^{2}}{90}T^{4},$ (2.1)
@@
for the case when $T\gg m$, where $m$ is the mass of the SF and $\lambda$ the self-interaction parameter between SF particles. Observations indicate that the self-interaction strength $\lambda$ must be extremely small but non-zero, as constrained by nucleosynthesis [50]. The mass $m$ of the SFDM is considered ultralight, which implies that any perturbation in its temperature is comparable to this mass in energy.
@@
Now, of particular interest for cosmology is the theoretical expectation that at high temperatures, symmetries that are spontaneously broken today were restored and, during the evolution of the Universe, there were phase transitions associated with the spontaneous breakdown of gauge symmetries [48]. So we expect that, as the Universe expanded, the SFDM particles and the other SM components, cooled down and underwent a phase transition at a critical temperature $T_{c}$ very early in the cosmic time. At this point, a spontaneous symmetry breaking (SSB) of the SF potential takes place.
@@
According to the standard cosmological model, after inflation, quantum fluctuations evolved into classical fluctuations, leading to the collapse of the SFDM and the formation of cosmic structures [15, 38, 31]. As the Universe expanded, the volume of these fluctuations increased, causing the SF to cool and form Bose-Einstein condensates. These condensates subsequently formed galaxy halos, where most SFDM particles settled into the ground state. However, gravitational forces caused the fluctuations to collapse again, decreasing their volume and increasing the halo’s temperature, resulting in some SFDM particles transitioning to excited states.
@@
In reference [39], it was demonstrated that if the majority of SFDM particles remain in the ground state while a portion of them occupy at least one excited state, the system stabilizes and remains stable for at least the age of the Universe, if the mass of the SFDM particle is sufficiently small, about $m\approx 10^{-22}\text{eV}/c^{2}$. This stability mechanism, involving a mixture of ground and excited states for SFDM particles within halos, is crucial for ensuring the formation and persistence of cosmic structures.
2.1 Previous Work and Unsolved Issues